### v7

@@ 3 Consumption and capital growth
Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Armenia 1997 - 2020 (20) 0.88 1.18 Aruba 1997 - 2001 (5) -0.22 8.44 Australia 1974 - 2019 (35) 0.97 1.08 Austria 1997 - 2021 (16) 0.90 1.13 Azerbaijan 1997 - 2020 (23) 0.25 2.20 Bahrain 2010 - 2013 (4) 0.59 1.32 Belgium 1997 - 2021 (19) 0.95 1.06 Bolivia 1999 - 2014 (12) 0.96 1.12 Botswana 1997 - 2000 (4) 0.76 1.27 Brazil 1998 - 2020 (22) 0.90 1.24 British Virgin Islands 1997 - 1998 (2) -1.24 -2.20 Bulgaria 1997 - 2017 (15) 1.03 1.19 Burkina Faso 2001 - 2019 (16) 0.96 1.28 Cameroon 1997 - 2019 (19) 0.82 2.71 Canada 1974 - 2022 (41) 0.91 1.08 Cape Verde 2009 - 2016 (8) 0.73 2.02 Chile 1998 - 2021 (22) 0.94 1.48 China 1993 - 2020 (23) 0.91 1.05 Colombia 1997 - 2022 (18) 1.03 2.39 Costa Rica 2014 - 2020 (6) 0.96 1.25 Croatia 1997 - 2021 (22) 0.81 2.09 Curaçao 2002 - 2016 (14) 0.59 0.51 Cyprus 1997 - 2021 (23) 1.06 1.16 Czechia 1995 - 2021 (19) 1.03 1.08 Côte d’Ivoire 1997 - 2000 (4) 1.03 1.05 Denmark 1997 - 2022 (22) 0.95 1.14 Dominican Republic 2007 - 2015 (9) 0.85 1.41 Ecuador 2009 - 2020 (12) 0.83 1.91 Egypt 1998 - 2015 (16) 0.87 2.14 El Salvador 2016 - 2019 (4) 0.02 -2.32 Estonia 1997 - 2021 (19) 0.94 1.11 Finland 1997 - 2022 (21) 1.02 1.20 France 1972 - 2021 (33) 0.91 1.16 Germany 1972 - 2022 (31) 0.95 1.14 Greece 1997 - 2021 (24) 0.76 1.47 Guatemala 2003 - 2021 (11) 0.85 1.32 Guinea 2005 - 2010 (6) 0.92 1.23 Honduras 2003 - 2015 (10) 0.60 0.85 Hong Kong 1997 - 2011 (13) 0.96 1.25 Hungary 1997 - 2021 (21) 0.93 1.10 Iceland 2002 - 2014 (12) 0.90 1.12 India 2000 - 2019 (14) 0.81 0.99 Indonesia 2019 - 2019 (1) 1.41 0.98 Iran 1998 - 2018 (19) -0.09 1.09 Ireland 1997 - 2021 (23) 0.98 1.23 Israel 1997 - 2017 (18) 0.93 1.72 Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Italy 1981 - 2021 (22) 0.94 1.12 Japan 1981 - 2021 (30) 0.95 1.13 Kazakhstan 1997 - 2022 (23) 0.92 1.16 Kuwait 2005 - 2017 (11) 0.96 1.13 Kyrgyzstan 1998 - 2021 (22) 0.77 1.29 Latvia 1998 - 2021 (20) 0.99 1.15 Lithuania 1997 - 2021 (18) 0.90 1.19 Luxembourg 1997 - 2021 (25) 1.00 1.02 Malaysia 2008 - 2015 (7) 0.99 1.19 Malta 1997 - 2021 (23) 0.92 1.04 Mauritius 2015 - 2018 (4) 0.98 0.99 Mexico 1997 - 2021 (22) 0.94 1.16 Moldova 1997 - 2019 (21) 0.89 1.27 Mongolia 2008 - 2021 (11) 0.90 1.06 Morocco 2000 - 2021 (17) 0.81 1.27 Netherlands 1997 - 2021 (20) 0.97 1.08 New Zealand 1997 - 2019 (18) 0.86 1.02 Nicaragua 2008 - 2019 (12) 0.94 1.13 Niger 1997 - 2019 (19) 0.78 1.87 Norway 1982 - 2021 (34) 0.88 1.08 Peru 2009 - 2021 (11) 0.96 1.30 Philippines 1997 - 2022 (20) 0.63 1.20 Poland 1997 - 2021 (15) 0.80 1.07 Portugal 1997 - 2022 (17) 0.98 1.08 Qatar 2004 - 2018 (12) 0.71 0.86 Romania 1997 - 2020 (14) 0.98 1.16 Russia 1997 - 2018 (14) 0.91 1.11 Saudi Arabia 2005 - 2009 (3) -1.23 0.02 Senegal 2018 - 2021 (3) 0.81 2.31 Serbia 1999 - 2021 (22) 0.78 1.67 Slovakia 1997 - 2021 (18) 0.91 1.18 Slovenia 1997 - 2021 (20) 0.90 1.07 South Africa 1997 - 2022 (21) 0.90 1.13 South Korea 1997 - 2020 (15) 0.96 1.05 Spain 1996 - 2021 (23) 0.96 1.11 Sweden 1974 - 2022 (41) 0.98 1.15 Switzerland 1993 - 2020 (24) 0.92 1.09 Tunisia 1997 - 2011 (12) 0.68 1.17 Turkey 2011 - 2015 (5) 1.16 1.54 USA 1972 - 2021 (44) 0.98 1.13 Ukraine 1997 - 2020 (23) 0.95 1.18 United Kingdom 1972 - 2021 (42) 0.98 1.21 Uruguay 2017 - 2017 (1) 0.44 0.98 Uzbekistan 2012 - 2021 (9) 0.95 1.14 Vanuatu 2003 - 2007 (3) 0.85 1.19 Venezuela 1999 - 2019 (20) 0.69 1.13
Note: Thrift theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 0$. Free growth theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 1$.
Table 5: Regression of $-\Delta c^{*}$ and $\Delta q_{c}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\text{ per thrift theory: }\Delta g(K)\cong-\Delta c^{*}\&\Delta q_{c}\cong 0$
$-\Delta c^{*}$ $\Delta q_{c}$
Regression of value shown on $\Delta g(K)$0.3760∗∗∗ 1.376∗∗∗
(0.1055) (0.1055)
Observations 1,548 1,548
R2 0.53993 0.88294
Within R2 0.31853 0.86226
Year fixed effects $\checkmark$ $\checkmark$
Country fixed effects $\checkmark$ $\checkmark$
4 Differences in test results
Fig. 2 and Table 3 show $\theta_{s}$ to average 0.064, and regression of $\Delta s^{*}$ on $\Delta g(K)$ to average $0.100$, across all years and countries GDP-weighted. Fig. 3 and Table 5, meanwhile, showed these overall averages as $-0.142$ for $\theta_{c}$, and as $-0.376$ for regression of $-\Delta c^{*}$ on $\Delta g(K)$. These differences might be explained by systematic mismeasurements of $C$, $K$ and $S_{net}$ in national accounts. In that case, the evidence would leave open the possibility of some role for net saving (6.4%) in capital growth.
There is another possible interpretation. Regression or correlation or covariance among variables such as net saving and capital growth suggests causal relationship among them, but does not in itself reveal the direction of causality. Thrift theory accepts the natural inference that net saving is the cause, and capital growth the effect. Logic and experience, however, also suggest that adverse shocks such as war or depression might lower capital values despite all saving and investment, and might then motivate negative net saving (dissaving), meaning suspension of replacement saving, in order to maintain consumption. When peace and prosperity return, higher incomes will allow restoration of replacement saving to offset depreciation. Capital growth need not depend on positive net saving at any point, but rather allows an upward normalization of net saving from a negative value to zero. In this interpretation, change in capital growth becomes the cause, rather than the effect, of change in net saving.
This possibility could help explain differences between the measurements of GDP-weighted $\varphi_{s}(0.936)$ and $\varphi_{c}(1.142)$, without need to assume systematic mismeasurements of $S_{net}$, $K$ and $C$. Note here that the finding $\varphi>1$ is possible in principle when gain in productivity allows growth in capital and consumption at the same time (see the quote from Mill in Appendix C.) We welcome further study to resolve these uncertainties.
@@ 5 Mechanics of free growth
The solution, we suggest, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.888i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
@@ 6 Optimum investment policy
Data and arguments adduced, with uncertainties discussed in Section 4, suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more. That would not mean book depreciation, as this study has stressed differences between book and market values. Up to a point, it should be possible to analyze the composition of market capital, and to model depreciation of the whole. A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to trust the market to maximize rate of return, and to sense the point where glut begins and returns fall. [^9] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator). Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption. Findings in this paper suggest review of such policies. These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains. Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
@@ 7 Inferences for national accounts and macroeconomics
Eq. (16) was presented as a tautology or logical certainty. Cash flow $F$ from all physical capital $K$ together at the collective scale, that is, means outflow of value from $K$ net of any inflow into $K$ from any source, meaning from pay $\Pi$ as well as plowback of cash flow. Thus the sum $F+\Pi$ is spent on consumption $C$ alone by exhaust of alternatives. Meanwhile cash flow is net profit $P$ less capital growth $\Delta K$ [^10] Or equivalently gross profit less capital growth less depreciation. This reasoning gave $F+\Pi=C$, $F=P-\Delta K$, and consequently $\Pi+P=C+\Delta K$ (Eq. (16)).
Kuznets and the other founders of national accounts, under the influence of Keynes, seem to have reasoned alike. National accounts are double-entry $T$ accounts where net product $C+\Delta K$ is equated to net "factor payments" $\Pi+P$. National accounts, however, equate capital growth to net saving/investment, and effectively write Eq. (16) as $\Pi+P=C+S_{net}$. We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth. If our argument proves correct, then national accounts, and the teachings of macroeconomics (macro) which they express, are constructed on a fundamental misperception. We would therefore welcome any critiques of our findings as a matter of academic and public interest.
@@ 11 Data sources
All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 86 of those countries which report all three of the factors, namely net saving, consumption and market-value capital, needed for deriving the thrift and free growth indexes. WID’s source for these data is national accounts.
Consumption $C$ in our text and equations is reproduced from the sum of Government Final Consumption Expenditure (mcongo) [^11] WID code and Private Expenditures of Households and NPISH (mconhn). This sums government expenditure GCE and personal consumption expenditure PCE. Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively. GDP, which we use only for weighting purposes in Figs. 1 and 2, is reproduced from GDP (mgdpro).
@@ 13 Displays
Eqs. (12) and (25) show $\theta_{s}+\varphi_{s}=1$ and $\theta_{c}+\varphi_{c}=1$. Some displays here and in the web appendix save space by showing $\varphi_{s}$ and $\varphi_{c}$ only, leaving $\theta_{s}$ and $\theta_{c}$ implicit as their complements to unity. The web appendix includes displays of $\varphi_{s}$ and $\varphi_{c}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\varphi_{s}$ and $\varphi_{c}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then .01, then .025, then continuing upward in increments of .025 to a maximum screen of .15. $\varphi_{s}$ and $\varphi_{c}$ are plotted for each country unscreened and at each of the seven successive levels of screening. All displays applied a screen of .01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 15 Discussion and conclusions
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence showing $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 1$ suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.

### v8

@@ 2 Net saving, capital growth and capital acceleration
Armenia 1997 - 2020 (20) 0.12 Aruba 1997 - 2001 (5) 1.22 Australia 1974 - 2019 (35) 0.03 Austria 1997 - 2021 (16) 0.10 Azerbaijan 1997 - 2020 (23) 0.75 Bahrain 2010 - 2013 (4) 0.41 Belgium 1997 - 2021 (19) 0.05 Bolivia 1999 - 2014 (12) 0.04 Botswana 1997 - 2000 (4) 0.24 Brazil 1998 - 2020 (22) 0.10 British Virgin Islands 1997 - 1998 (2) 2.24 Bulgaria 1997 - 2017 (15) -0.03 Burkina Faso 2001 - 2019 (16) 0.04 Cameroon 1997 - 2019 (19) 0.18 Canada 1974 - 2022 (41) 0.09 Cape Verde 2009 - 2016 (8) 0.27 Chile 1998 - 2021 (22) 0.06 China 1981 - 2020 (33) 0.07 Colombia 1997 - 2022 (18) -0.03 Costa Rica 2014 - 2020 (6) 0.04 Croatia 1997 - 2021 (22) 0.19 Curaçao 2002 - 2016 (14) 0.41 Cyprus 1997 - 2021 (23) -0.06 Czechia 1995 - 2021 (19) -0.03 Côte d’Ivoire 1997 - 2000 (4) -0.03 Denmark 1997 - 2022 (22) 0.05 Dominican Republic 2007 - 2015 (9) 0.15 Ecuador 2009 - 2020 (12) 0.17 Egypt 1998 - 2015 (16) 0.13 El Salvador 2016 - 2019 (4) 0.98 Estonia 1997 - 2021 (19) 0.06 Finland 1997 - 2022 (21) -0.02 France 1972 - 2021 (33) 0.09 Germany 1972 - 2022 (31) 0.05 Greece 1997 - 2021 (24) 0.24 Guatemala 2003 - 2021 (11) 0.15 Guinea 2005 - 2010 (6) 0.08 Honduras 2003 - 2015 (10) 0.40 Hong Kong 1997 - 2011 (13) 0.04 Hungary 1997 - 2021 (21) 0.07 Iceland 2002 - 2014 (12) 0.10 India 1997 - 2019 (16) 0.14 Indonesia 2019 - 2019 (1) -0.41 Iran 1998 - 2018 (19) 1.09 Ireland 1997 - 2021 (23) 0.02 Israel 1997 - 2017 (18) 0.07 Italy 1972 - 2021 (29) 0.05 Japan 1972 - 2021 (38) 0.05 Kazakhstan 1997 - 2022 (23) 0.08 Kuwait 2005 - 2017 (11) 0.04 Kyrgyzstan 1998 - 2021 (22) 0.23 Latvia 1998 - 2021 (20) 0.01 Lithuania 1997 - 2021 (18) 0.10 Luxembourg 1997 - 2021 (25) 0.00 Malaysia 2008 - 2015 (7) 0.01 Malta 1997 - 2021 (23) 0.08 Mauritius 2015 - 2018 (4) 0.02 Mexico 1997 - 2021 (22) 0.06 Moldova 1997 - 2019 (21) 0.11 Mongolia 2008 - 2021 (11) 0.10 Morocco 2000 - 2021 (17) 0.19 Netherlands 1997 - 2021 (20) 0.03 New Zealand 1997 - 2019 (18) 0.14 Nicaragua 2008 - 2019 (12) 0.06 Niger 1997 - 2019 (19) 0.22 Norway 1982 - 2021 (34) 0.12 Peru 2009 - 2021 (11) 0.04 Philippines 1997 - 2022 (20) 0.37 Poland 1997 - 2021 (15) 0.20 Portugal 1997 - 2022 (17) 0.02 Qatar 2004 - 2018 (12) 0.29 Romania 1997 - 2020 (14) 0.02 Russia 1997 - 2018 (14) 0.09 Saudi Arabia 2005 - 2009 (3) 2.23 Senegal 2018 - 2021 (3) 0.19 Serbia 1999 - 2021 (22) 0.22 Slovakia 1997 - 2021 (18) 0.09 Slovenia 1997 - 2021 (20) 0.10 South Africa 1997 - 2022 (21) 0.10 South Korea 1997 - 2020 (15) 0.04 Spain 1972 - 2021 (42) 0.06 Sweden 1974 - 2022 (41) 0.02 Switzerland 1993 - 2020 (24) 0.08 Tunisia 1997 - 2011 (12) 0.32 Turkey 2011 - 2015 (5) -0.16 USA 1972 - 2021 (44) 0.02 Ukraine 1997 - 2020 (23) 0.05 United Kingdom 1972 - 2021 (42) 0.02 Uruguay 2017 - 2017 (1) 0.56 Uzbekistan 2012 - 2021 (9) 0.05 Vanuatu 2003 - 2007 (3) 0.15 Venezuela 1999 - 2019 (20) 0.31
Note: Thrift theory predicts $\theta_{s}^{*}=1$.
Fig. 2: Average $\theta_{s}^{*}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
3 Interpretation of test results
The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, and does not of itself reveal causal relatedness between those two variables. Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
$\theta_{s}$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $\theta_{s}^{*}=0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained. Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$. Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation. This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption. Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income. This interpretation could fit the appearance in Fig. 2 that $\frac{\Delta s^{*}}{\Delta g(K)}$, shown as $\theta^{*}_{s}$, found higher values during the period of the dot com and subprime crises.
In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)). Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\frac{\Delta s^{*}}{\Delta g(K)}$ as does thrift theory, but from an opposite causality.
Although these results cannot exclude the possibility that some net saving has effect, data are easiest to reconcile with free growth theory, where all capital growth is explained by productivity gain. Pending further study, we will proceed on this interpretation.
@@ 4 Mechanics of free growth
The solution, we suggest, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
@@ 5 Optimum investment policy
Data and arguments suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more. That would not mean book depreciation, as this study has stressed differences between book and market values. Up to a point, it should be possible to analyze the composition of market capital, and to model depreciation of the whole. A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to trust the market to maximize rate of return, and to sense the point where glut begins and returns fall. [^6] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator). Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption. Findings in this paper suggest review of such policies. These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains. Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
@@ 6 Inferences for national accounts and macroeconomics
We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth. If our argument proves correct, then national accounts, and the teachings of macroeconomics (macro) which they express, are constructed on a fundamental misperception. We would therefore welcome any critiques of our findings as a matter of academic and public interest.
@@ 10 Data sources
All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 92 of those countries which report both factors, net saving and market-value capital, needed for deriving the thrift indexes. WID’s source for these data is national accounts.
Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively. GDP, which we use only for weighting purposes in Figs. 1 and 2, is reproduced from GDP (mgdpro).
@@ 12 Displays
The web appendix includes displays of $\theta_{s}^{*}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\theta_{s}^{*}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15. $\theta_{s}^{*}$ is plotted for each country unscreened and at each of the seven successive levels of screening. All displays applied a screen of 0.01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 14 Discussion and conclusions
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.